OpenStudy (anonymous):

A rectangular field that is 119 yards in area is to be covered with sod. It is kow that the field is 3 yards longer than twice the width Find the dimensions of the field.

OpenStudy (chihiroasleaf):

the field is 3 yards longer than twice the width , it means l = 2w + 3, the area of the yards is l x w, so you have l x w = 119 (2w+3) w = 119 simplify and solve this equation, then you'll get w

OpenStudy (anonymous):

how do i solve?

OpenStudy (chihiroasleaf):

use distributive property, then simplify the equation, can you?

OpenStudy (anonymous):

i come up with the wrong answer :(

OpenStudy (anonymous):

i came up with 116 = 3x,,, the 38,6666

OpenStudy (chihiroasleaf):

(2w+3) w = 119 2w^2 + 3w = 119 2w^2 + 3w - 119 = 0 now, you have quadratic equation now, solve the equation

OpenStudy (anonymous):

oh,,, ii was there once,, but i thought it was wrong

OpenStudy (anonymous):

so i have to use the quadratic formula?

OpenStudy (chihiroasleaf):

yes.., if it cannot be factorized